OpenStudy (anonymous):

Suppose E(x)=5 and E[x(x-1)]=27.5. What is E[x^2]?

OpenStudy (perl):

you can use properties of expected value

OpenStudy (perl):

Properties of expectation Linearity When a is constant and X,Y are random variables: E(aX) = aE(X) E(X+Y) = E(X) + E(Y) Constant When c is constant: E(c) = c Product When X and Y are independent random variables: E(X *Y) = E(X) * E(Y)

OpenStudy (perl):

we can use this as a guide

OpenStudy (perl):

E(x)=5 and E[x(x-1)] = 27.5 but E[x(x-1)] =E(x^2 - x)= E(x^2) - E(x)

OpenStudy (perl):

now substitute

OpenStudy (perl):

E(x^2) - E(x) = 27.5 E(x^2) = 27.5 + E(x) E(x^2) = 27.5 + 5